A pyramid has a square base. Find the surface area of the pyramid if its slant height is 10 centimeters and the sides of its base are 5 centimeters long.
Answer by checkley71(8403)

A SQUARE PYRAMID HAS 4 TRIANGLE SIDES WITH A BASE OF 5 CM & A HEIGHT OF 10CM.
AREA OF EACH TRIANGLE=(B*H)/2
THUS WE HAVE 4(5*10)/2=4*50/2=200/2=100 SQUARE CM.